# Association Scheme on Triples from the Unitary Group

### v1

@@ Abstract
An association scheme on triples (AST) is a three-dimensional analogue of a classical association scheme. Similar to how a transitive group action produces a Schurian classical association scheme, a two-transitive group action produces an AST. This paper describes the relations and intersection numbers of ASTs from the finite unitary groups acting on the respective isotropic lines.
@@ 1 Introduction
A classical association scheme on a set $\Omega$ is a partition of $\Omega\times\Omega$ satisfying certain regularity properties. By providing a single framework with which to study several discrete structures, classical association schemes have proven useful in the study of combinatorial objects such as graphs, codes, and designs.
Introduced in [5] as a higher-dimensional analogue of a classical association scheme, an association scheme on triples (AST) on a set $\Omega$ is a partition of $\Omega\times\Omega\times\Omega$ satisfying certain symmetry requirements. Similar to how classical association schemes arise from transitive group actions, ASTs arise naturally from two-transitive group actions. The parameters of several families of such ASTs were obtained in [1].
However, the parameters of some families of such ASTs remain undetermined. Indeed, due to the more involved action of the finite unitary groups on isotropic lines, the parameters of the corresponding ASTs were left incomplete. In this paper, we continue the work in [1] by leveraging the matrix representation of the unitary groups to describe explicitly the relations and intersection numbers of the ASTs these groups produce. The parameters are expressed in the following main result.
@@ 9 $p_{i44}^{\ell}=p_{4j4}^{\ell}=p_{44k}^{\ell}=p_{444}^{\ell}=0$
This is equivalent to the existence of $\zeta]=[bu+v+cw]$ in $([b_{k}u+v+c_{k}w])^{U(3,q^{2})_{u,v}}$ satisfying the following conditions.
@@
It remains to show that if $p_{ijk}^{4}\neq 0$, then there are exactly $q+1$ elements $\zeta]$ in the orbit $([b_{k}u+v+c_{k}w])^{U(3,q^{2})_{u,v}}$ such that Equations (9) and (10) are true. To see this, fix $\zeta]=[bu+v+cw]\in\Gamma_{ijk}^{4}([u],[v],[au+v]).$ Recall that the elements of $([b_{k}u+v+c_{k}w])^{U(3,q^{2})_{u,v}}$ are $[du+v+ew]$, where $d\in bF_{0}^{\times}$, and $e\in F^{\times}$ satisfies $d+\overline{d}+e\overline{e}=0$. For each of the $q+1$ elements $e\in F^{\times}$ satisfying $e\overline{e}=c\overline{c}$, the element $[bu+v+ew]$ satisfies Equations (9) and (10). This accounts for $q+1$ elements of $\Gamma_{ijk}^{4}([u],[v],[au+v])$. To complete the proof, we show that $du+v+ew]\notin\Gamma_{ijk}^{4}([u],[v],[au+v])$ if $b\neq d$. Indeed, let $d=by$ for some $1\neq y\in F_{0}^{\times}$. If $du+v+ew]\in\Gamma_{ijk}^{4}([u],[v],[au+v])$, then Equation (10) yields $b_{i}d(\overline{d}+a)a^{-1}\in F_{0}^{\times}$. Since $b_{i}b(\overline{b}+a)a^{-1}$ is also in $F_{0}^{\times}$, we obtain
@@
whenever $k,\ell\in\{5,\ldots,m\}$ are distinct.
@@
Multiplying the corresponding sides of the above equations yields $h^{2}+he\overline{e}=\overline{h}^{2}+\overline{h}e\overline{e}$. Since $h^{2}=\overline{h}^{2}$, it follows that $h=\overline{h}$. That is, $h\in F_{0}^{\times}$. Since $\operatorname{char}F=2$ and $k\geq 5$, we obtain a contradiction. ∎

### v2

@@ Abstract
An association scheme on triples (AST) is a three-dimensional analogue of a classical association scheme. Since much of the algebraic and combinatorial properties of ASTs remain unknown, it is natural to determine the structure constants of ASTs that are constructed similarly to classical counterparts. Similar to how a transitive group action produces a Schurian classical association scheme, a two-transitive group action produces an AST. The parameters of several such ASTs are known; however, the intersection numbers of the ASTs from the finite unitary group remained undetermined. In this paper, the parameters of the AST from the finite unitary group are obtained through explicit descriptions of the AST’s relations and certain equations on the underlying field. In particular, we observe that an adjacency hypermatrix of this AST generates a one-dimensional ternary algebra.
@@ 1 Introduction
A classical association scheme on a set $\Omega$ is a partition of $\Omega\times\Omega$ satisfying certain regularity properties [4]. By providing a single framework with which to study several discrete structures, classical association schemes have proven useful in the study of combinatorial objects such as graphs, codes, and designs.
Introduced in [8] as a higher-dimensional analogue of a classical association scheme, an association scheme on triples (AST) on a set $\Omega$ is a partition of $\Omega\times\Omega\times\Omega$ satisfying certain symmetry requirements. In that paper, the authors demonstrated several relationships between ASTs and known combinatorial structures, including classical association schemes, 2-designs, and two-graphs. Despite being introduced in 1990 and the various avenues open to exploration, the area is still in its nascency, with only a few papers published since its conception. Several such papers appeared recently, including [5], [1], [3], and [2].
Still, little is known about the algebraic or combinatorial structure of ASTs. To this end, it is natural to consider the structural parameters of families of ASTs that admit well-understood automorphism groups. One such approach was taken in [5], where ASTs under consideration were invariant under some transitive cyclic subgroup. In this paper, we instead continue an approach initiated in [8], and furthered in [3]. Similar to how classical association schemes arise from transitive group actions, [8] shows that ASTs arise naturally from two-transitive group actions. As there are few major families of two-transitive groups (as listed in [6]), and since this situation mirrors the familiar classical case, it is a natural goal to determine the structural constants of these ASTs. The parameters of several families of such ASTs were obtained in [3]. Moreover, among the ASTs from the broadest classes of two-transitive groups, only the parameters of the ASTs from the Suzuki, Ree, and projective unitary groups remain undetermined. We continue the work in [3] by leveraging the matrix representation of the unitary groups to describe explicitly the relations and intersection numbers of the ASTs these groups produce. In order to afford convenient descriptions of the AST parameters, we differ slightly from the approach to the computations that was taken in [3]. The initial approach is similar, leveraging a convenient basis for the unitary space and the projective transformations. However, our method diverges when we phrase the necessary and sufficient conditions for an intersection number to be nonzero in terms of certain inclusions in the fixed field. This yields the intersection numbers described in Theorem 3.2. In particular, we find that a one-dimensional ternary subalgebra is generated by a particular adjacency hypermatrix.
@@ 9 $p_{i44}^{\ell}=p_{4j4}^{\ell}=p_{44k}^{\ell}=p_{444}^{\ell}=0$
This is equivalent to the existence of $\zeta]=[bu+v+cw]$ in $([b_{k}u+v+c_{k}w])^{U(3,q^{2})_{[u],[v]}}$ satisfying the following conditions.
@@
It remains to show that if $p_{ijk}^{4}\neq 0$, then there are exactly $q+1$ elements $\zeta]$ in the orbit $([b_{k}u+v+c_{k}w])^{U(3,q^{2})_{[u],[v]}}$ such that Equations (9) and (10) are true. To see this, fix $\zeta]=[bu+v+cw]\in\Gamma_{ijk}^{4}([u],[v],[au+v]).$ Recall that the elements of $([b_{k}u+v+c_{k}w])^{U(3,q^{2})_{[u],[v]}}$ are $[du+v+ew]$, where $d\in bF_{0}^{\times}$, and $e\in F^{\times}$ satisfies $d+\overline{d}+e\overline{e}=0$. For each of the $q+1$ elements $e\in F^{\times}$ satisfying $e\overline{e}=c\overline{c}$, the element $[bu+v+ew]$ satisfies Equations (9) and (10). This accounts for $q+1$ elements of $\Gamma_{ijk}^{4}([u],[v],[au+v])$. To complete the proof, we show that $du+v+ew]\notin\Gamma_{ijk}^{4}([u],[v],[au+v])$ if $b\neq d$. Indeed, let $d=by$ for some $1\neq y\in F_{0}^{\times}$. If $du+v+ew]\in\Gamma_{ijk}^{4}([u],[v],[au+v])$, then Equation (10) yields $b_{i}d(\overline{d}+a)a^{-1}\in F_{0}^{\times}$. Since $b_{i}b(\overline{b}+a)a^{-1}$ is also in $F_{0}^{\times}$, we obtain
@@
whenever $k,\ell\in\{5,\ldots,m\}$ are distinct. Moreover, if $5\leq i\leq q+4$ then $\operatorname{Span}\left\{A_{i}\right\}$ cannot be a nontrivial subalgebra of $\operatorname{Span}\left\{A_{4},\ldots A_{q+4}\right\}$.
@@
Multiplying the corresponding sides of the above equations yields $h^{2}+he\overline{e}=\overline{h}^{2}+\overline{h}e\overline{e}$. Since $h^{2}=\overline{h}^{2}$, it follows that $h=\overline{h}$. That is, $h\in F_{0}^{\times}$. Since $\operatorname{char}F=2$ and $k\geq 5$, we obtain a contradiction.
The final statement holds since $p_{iii}^{i}=0$ for $5\leq i\leq q+4$